OpenStudy (anonymous):

In the right triangle, find the length of the side not given. Give an exact answer and an approximation to three decimal places. a=11, b=11 What is the exact value of c?

OpenStudy (saifoo.khan):

is c the hypotenuse?

jimthompson5910 (jim_thompson5910):

a^2+b^2 = c^2 11^2+11^2 = c^2 121 + 121 = c^2 242 = c^2 c^2 = 242 c = sqrt(242) c = 11sqrt(2) So \[\Large c = 11\sqrt{2}\]